OpenStudy (anonymous):

Read the following question carefully and relate your answer to the number of protons, electrons, and neutrons in an atom. Please do not spell out the number. The mass number of a certain atom is 35. Its atomic number is 17. How many neutrons will this atom have?

OpenStudy (lanre):

mass number=proton +neutron= 35. atomic number= proton or electron=17 Neutron=35-17=18 Therefore Neutron=18

OpenStudy (anonymous):

how did u get that i have lots like this one

OpenStudy (lanre):

see the steps. Mass number of any atom=proton +neutron. The atomic number of any atom is the same as electron or proton number of that atom. Now, the mass number was given as 35. So, Neutron will be 35 - proton, where proton is the same as atomic number, which is given as 17. So, remove 17 FROM 35 = 18

OpenStudy (anonymous):

Ok, so (An atom has 12 electrons and a mass number of 24. How many neutrons does it have?) 24-12= 12?

OpenStudy (lanre):

yes

OpenStudy (anonymous):

well how would i solve this one? If an atom has 16 protons in its nucleus, how many electrons must it have?

OpenStudy (anonymous):

A Proton is positive and an electron is negative. For an atom to be neutral, the charge must add up to 0. So the number of electrons in your case is 16, unless the atom is an ion.

OpenStudy (anonymous):

thans :)

OpenStudy (anonymous):

You're welcome. Glad I could help.